OpenStudy (anonymous):

Identify the number as rational or irrational. Explain. 291.87

OpenStudy (amoodarya):

it is rational because we can write it as 291.87 =29187/100 may be simplified !

OpenStudy (anonymous):

thank you

OpenStudy (amoodarya):

in this case does not simplified but 3.5 =35/10 =7/2